OpenStudy (anonymous):

Will is 8 years old Jan is 22 years old In how many years time will Jan be twice as old as Will? In.............years time Explain how you figured it out.

OpenStudy (exoexoexo):

use the equation 2(x+8)=(x+22)

OpenStudy (exoexoexo):

x is the amount of time that will's age will be half of Jan's age

OpenStudy (anonymous):

is it x = 6

OpenStudy (anonymous):

so then it would be 12 years in time? ..........no......come........back...:'(

OpenStudy (irishboy123):

\(W_o = 8\) \(J_o = 22\) \(W(t) = W_o + t \\ = 8 + t\) \(J(t) = J_o + t \\ = 22 + t\) \(J(t) = 2 W(t) \\ \implies 22 + t = 2( 8 + t) \\ \implies t = 6\) check \(W(6) = 14\) \(J(6) = 28\) \(J =2 \times W\) \(\Large \checkmark\)

OpenStudy (exoexoexo):

\[2(x+8)=(x+22)\] \[2x+16=x=22\] \[2x+16-16=x+22-16\] \[2x=x+6\] \[2x-x=x-x+6\] \[x=6\] so the answer is 6 years

OpenStudy (anonymous):

ok THXXXX!